What is the present value of a property with next year's NOI of $100,000, which will grow at 3% per year and discounted at a rate of 8%? Use PV=FV/e^Tk
Answer #1

NOI=   $100,000  
growth rate=   3%  
k= Required or Discount rate=   8%  
t is year that is annual. so t=1   $1  
so e^tk= e^8%*1= e^0.08      
      
e^0.08= ((i)^1/1) + ((i)^2/(2*1))+ ((i)^3/(3*2*1)) + ((i)^4/(4*3*2*1))      
e^0.08 = ((0.08)^1/1) + ((0.08)^2/(2*1))+ ((0.08)^3/(3*2*1)) + ((0.08)^4/(4*3*2*1))      
0.08328704      
      
Present value = NOI/(e^tk - g)      
100000/(   0.08328704   -0.03)
$1,876,629

So Present value of property is $1,876,629

Please thumbs up or post comments if any query